Calculate the acceleration of the electron which electronic charge is given as(1.60×10^-19c) and mass of electron as (9.1×10^-3kg)


1
Expert's answer
2021-09-03T15:25:59-0400

Suppose the electric field in the place of electron's location be "E". Then the force on election is:


"F =eE"

where "e = 1.6\times 10^{-19}C" is the charge of electron. According to the second Newton's law, the acceleration of the electron is:


"a = \dfrac{F}{m} = \dfrac{eE}{m}"

where "m = 9.1\times 10^{-31}kg" is the mass of electron. Substituting the particelar value of "E" one can find the particular value of acceleration.


Answer. "a = \dfrac{eE}{m}".